# TimeFilter: Patch-Specific Spatial-Temporal Graph Filtration for Time Series Forecasting

### v1

@@ B.2 Long-term Forecasting
ETTm1 96 0.313 0.354 0.334 0.368 0.319 0.359 0.319 0.366 0.325 0.361 0.335 0.373 0.329 0.367 0.404 0.426 0.338 0.375 0.346 0.374 0.379 0.419
192 0.356 0.380 0.377 0.391 0.369 0.383 0.376 0.397 0.375 0.389 0.372 0.390 0.367 0.385 0.450 0.451 0.374 0.387 0.382 0.391 0.426 0.441
336 0.386 0.402 0.426 0.420 0.394 0.402 0.417 0.422 0.405 0.412 0.403 0.411 0.399 0.410 0.532 0.515 0.410 0.411 0.415 0.415 0.445 0.459
720 0.452 0.437 0.491 0.459 0.460 0.442 0.481 0.458 0.466 0.447 0.461 0.442 0.454 0.439 0.666 0.589 0.478 0.450 0.473 0.451 0.543 0.490
Avg. 0.377 0.393 0.407 0.410 0.386 0.397 0.398 0.411 0.393 0.403 0.393 0.404 0.387 0.400 0.513 0.496 0.400 0.406 0.404 0.408 0.448 0.452
ETTm2 96 0.169 0.255 0.180 0.264 0.176 0.257 0.177 0.262 0.180 0.261 0.176 0.266 0.175 0.259 0.287 0.366 0.187 0.267 0.193 0.293 0.203 0.287
192 0.235 0.299 0.250 0.309 0.243 0.303 0.247 0.307 0.246 0.306 0.240 0.307 0.241 0.302 0.414 0.492 0.249 0.309 0.284 0.361 0.269 0.328
336 0.293 0.336 0.311 0.348 0.303 0.341 0.312 0.346 0.319 0.352 0.304 0.345 0.305 0.343 0.597 0.542 0.321 0.351 0.382 0.429 0.325 0.366
720 0.390 0.393 0.412 0.407 0.400 0.398 0.414 0.403 0.405 0.401 0.406 0.400 0.402 0.400 1.730 1.042 0.408 0.403 0.558 0.525 0.421 0.415
Avg. 0.272 0.321 0.288 0.332 0.281 0.325 0.288 0.330 0.287 0.330 0.282 0.330 0.281 0.326 0.757 0.610 0.291 0.333 0.354 0.402 0.305 0.349
ETTh1 96 0.370 0.394 0.386 0.405 0.377 0.394 0.390 0.411 0.381 0.399 0.382 0.398 0.414 0.419 0.423 0.448 0.384 0.402 0.397 0.412 0.376 0.419
192 0.413 0.420 0.441 0.436 0.424 0.422 0.442 0.442 0.435 0.431 0.427 0.425 0.460 0.445 0.471 0.474 0.436 0.429 0.446 0.441 0.420 0.448
336 0.450 0.440 0.487 0.458 0.459 0.442 0.480 0.468 0.480 0.452 0.465 0.445 0.501 0.466 0.570 0.546 0.491 0.469 0.489 0.467 0.459 0.465
720 0.448 0.457 0.503 0.491 0.463 0.459 0.494 0.488 0.499 0.488 0.472 0.468 0.500 0.488 0.653 0.621 0.521 0.500 0.513 0.510 0.506 0.507
Avg. 0.420 0.428 0.454 0.447 0.431 0.429 0.452 0.452 0.449 0.442 0.437 0.434 0.469 0.454 0.529 0.522 0.458 0.450 0.461 0.457 0.440 0.460
ETTh2 96 0.283 0.337 0.297 0.349 0.292 0.343 0.328 0.371 0.297 0.347 0.309 0.359 0.302 0.348 0.745 0.584 0.340 0.374 0.340 0.394 0.358 0.397
192 0.362 0.392 0.380 0.400 0.367 0.389 0.402 0.414 0.373 0.394 0.390 0.406 0.388 0.400 0.877 0.656 0.402 0.414 0.482 0.479 0.429 0.439
336 0.404 0.424 0.428 0.432 0.412 0.424 0.435 0.443 0.410 0.426 0.426 0.444 0.426 0.433 1.043 0.731 0.452 0.452 0.591 0.541 0.496 0.487
720 0.407 0.433 0.427 0.445 0.419 0.438 0.417 0.441 0.411 0.433 0.445 0.464 0.431 0.446 1.104 0.763 0.462 0.468 0.839 0.661 0.463 0.474
Avg. 0.364 0.397 0.383 0.407 0.373 0.399 0.396 0.417 0.385 0.408 0.373 0.400 0.387 0.407 0.942 0.684 0.414 0.427 0.563 0.519 0.437 0.449
Weather 96 0.153 0.199 0.174 0.214 0.156 0.202 0.163 0.212 0.166 0.208 0.159 0.218 0.177 0.218 0.158 0.230 0.172 0.220 0.195 0.252 0.217 0.296
192 0.202 0.246 0.221 0.254 0.207 0.250 0.212 0.254 0.217 0.253 0.211 0.266 0.225 0.259 0.206 0.277 0.219 0.261 0.237 0.295 0.276 0.336
336 0.260 0.289 0.278 0.296 0.262 0.291 0.272 0.299 0.282 0.300 0.267 0.310 0.278 0.297 0.272 0.335 0.280 0.306 0.282 0.331 0.339 0.380
720 0.342 0.341 0.358 0.349 0.343 0.343 0.350 0.348 0.356 0.351 0.352 0.362 0.354 0.348 0.398 0.418 0.365 0.359 0.345 0.382 0.403 0.428
Avg. 0.239 0.269 0.258 0.279 0.242 0.272 0.249 0.278 0.255 0.278 0.247 0.289 0.259 0.281 0.259 0.315 0.259 0.287 0.265 0.315 0.309 0.360
Electricity 96 0.133 0.230 0.148 0.240 0.141 0.235 0.165 0.274 0.143 0.233 0.173 0.275 0.195 0.285 0.219 0.314 0.168 0.272 0.210 0.302 0.193 0.308
192 0.154 0.248 0.162 0.253 0.159 0.252 0.184 0.292 0.158 0.248 0.195 0.288 0.199 0.289 0.231 0.322 0.184 0.289 0.210 0.305 0.201 0.315
336 0.162 0.261 0.178 0.269 0.173 0.268 0.195 0.302 0.178 0.269 0.206 0.300 0.215 0.305 0.246 0.337 0.198 0.300 0.223 0.319 0.214 0.329
720 0.184 0.284 0.225 0.317 0.201 0.295 0.231 0.332 0.218 0.305 0.231 0.335 0.256 0.337 0.280 0.363 0.220 0.320 0.258 0.350 0.246 0.355
Avg. 0.158 0.256 0.178 0.270 0.169 0.263 0.194 0.300 0.174 0.264 0.201 0.300 0.216 0.304 0.244 0.334 0.193 0.295 0.225 0.319 0.214 0.327
Traffic 96 0.375 0.251 0.395 0.268 0.426 0.276 0.605 0.344 0.376 0.251 0.570 0.310 0.544 0.359 0.522 0.290 0.593 0.321 0.650 0.396 0.587 0.366
192 0.395 0.262 0.417 0.276 0.458 0.289 0.613 0.359 0.398 0.261 0.577 0.321 0.540 0.354 0.530 0.293 0.617 0.336 0.598 0.370 0.604 0.373
336 0.414 0.271 0.433 0.283 0.486 0.297 0.642 0.376 0.415 0.269 0.588 0.324 0.551 0.358 0.558 0.305 0.629 0.336 0.605 0.373 0.621 0.383
720 0.445 0.289 0.467 0.302 0.498 0.313 0.702 0.401 0.447 0.287 0.597 0.337 0.586 0.375 0.589 0.328 0.640 0.350 0.645 0.394 0.626 0.382
Avg. 0.407 0.268 0.428 0.282 0.467 0.294 0.641 0.370 0.409 0.267 0.583 0.323 0.555 0.362 0.550 0.304 0.620 0.336 0.625 0.383 0.610 0.376
Solar-Energy 96 0.193 0.223 0.203 0.237 0.197 0.241 0.208 0.243 0.200 0.230 0.283 0.353 0.234 0.286 0.310 0.331 0.250 0.292 0.290 0.378 0.242 0.342
192 0.226 0.249 0.233 0.261 0.231 0.263 0.258 0.281 0.229 0.253 0.316 0.374 0.267 0.310 0.734 0.725 0.296 0.318 0.320 0.398 0.285 0.380
336 0.235 0.261 0.248 0.273 0.241 0.268 0.293 0.311 0.243 0.269 0.347 0.393 0.290 0.315 0.750 0.735 0.319 0.330 0.353 0.415 0.282 0.376
720 0.239 0.268 0.249 0.275 0.250 0.281 0.290 0.315 0.245 0.272 0.348 0.389 0.289 0.317 0.769 0.765 0.338 0.337 0.357 0.413 0.357 0.427
Avg. 0.223 0.250 0.233 0.262 0.230 0.264 0.262 0.288 0.229 0.256 0.324 0.377 0.270 0.307 0.641 0.639 0.301 0.319 0.330 0.401 0.292 0.381
@@ B.3 Short-term Forecasting
Tab. 10 present the full results for short-term forecasting. The look-back horizon $L$ is set to $96$ and the forecasting horizon $T\in\{12,24,48\}$. Across all four PEMS datasets, TimeFilter consistently delivers the highest performance, particularly with a significant improvement on PEMS08, showcasing its effectiveness and reliability.
Models TimeFilter iTransformer Leddam SOFTS PatchTST Crossformer TimesNet TiDE DLinear SCINet FEDformer
@@
PEMS03 12 0.063 0.165 0.071 0.174 0.068 0.174 0.064 0.165 0.099 0.216 0.090 0.203 0.085 0.192 0.178 0.305 0.122 0.243 0.066 0.172 0.126 0.251
24 0.079 0.185 0.093 0.201 0.094 0.202 0.083 0.188 0.142 0.259 0.121 0.240 0.118 0.223 0.257 0.371 0.201 0.317 0.085 0.198 0.149 0.275
48 0.110 0.222 0.125 0.236 0.140 0.254 0.114 0.223 0.211 0.319 0.202 0.317 0.155 0.260 0.379 0.463 0.333 0.425 0.127 0.238 0.227 0.348
Avg. 0.084 0.191 0.096 0.204 0.101 0.210 0.087 0.192 0.151 0.265 0.138 0.253 0.119 0.271 0.271 0.380 0.219 0.295 0.093 0.203 0.167 0.291
PEMS04 12 0.068 0.167 0.078 0.183 0.076 0.182 0.074 0.176 0.105 0.224 0.098 0.218 0.087 0.195 0.219 0.340 0.148 0.272 0.073 0.177 0.138 0.262
24 0.080 0.183 0.095 0.205 0.097 0.209 0.088 0.194 0.153 0.257 0.131 0.256 0.103 0.215 0.292 0.398 0.224 0.340 0.084 0.193 0.177 0.293
48 0.101 0.209 0.120 0.233 0.132 0.249 0.110 0.219 0.229 0.339 0.205 0.326 0.136 0.250 0.409 0.478 0.335 0.437 0.099 0.211 0.270 0.368
Avg. 0.083 0.186 0.098 0.207 0.102 0.213 0.091 0.196 0.162 0.273 0.145 0.267 0.109 0.220 0.307 0.405 0.236 0.350 0.085 0.194 0.195 0.308
PEMS07 12 0.055 0.150 0.067 0.165 0.066 0.164 0.057 0.152 0.095 0.207 0.094 0.200 0.082 0.181 0.173 0.304 0.115 0.242 0.068 0.171 0.109 0.225
24 0.068 0.166 0.088 0.190 0.079 0.185 0.073 0.173 0.150 0.262 0.139 0.247 0.101 0.204 0.271 0.383 0.210 0.329 0.119 0.225 0.125 0.244
48 0.089 0.193 0.110 0.215 0.115 0.228 0.096 0.195 0.253 0.340 0.311 0.369 0.134 0.238 0.446 0.495 0.398 0.458 0.149 0.237 0.165 0.288
Avg. 0.071 0.170 0.088 0.190 0.087 0.192 0.075 0.173 0.166 0.270 0.181 0.272 0.106 0.208 0.297 0.394 0.241 0.343 0.112 0.211 0.133 0.282
PEMS08 12 0.064 0.162 0.079 0.182 0.070 0.173 0.074 0.171 0.168 0.232 0.165 0.214 0.112 0.212 0.227 0.343 0.154 0.276 0.087 0.184 0.173 0.273
24 0.079 0.182 0.115 0.219 0.091 0.200 0.104 0.201 0.224 0.281 0.215 0.260 0.141 0.238 0.318 0.409 0.248 0.353 0.122 0.221 0.210 0.310
48 0.105 0.214 0.186 0.235 0.145 0.261 0.164 0.253 0.321 0.354 0.315 0.335 0.198 0.283 0.497 0.510 0.440 0.470 0.189 0.270 0.320 0.394
Avg. 0.083 0.186 0.127 0.212 0.102 0.211 0.114 0.208 0.238 0.289 0.232 0.270 0.150 0.244 0.347 0.421 0.281 0.366 0.133 0.225 0.234 0.326

### v2

@@ B.2 Long-term Forecasting
MSE
MAE
ETTm1 96 0.313 0.354 0.324 0.354 0.334 0.368 0.319 0.359 0.319 0.366 0.325 0.361 0.335 0.373 0.329 0.367 0.404 0.426 0.338 0.375 0.346 0.374 0.379 0.419
192 0.356 0.380 0.369 0.379 0.377 0.391 0.369 0.383 0.376 0.397 0.375 0.389 0.372 0.390 0.367 0.385 0.450 0.451 0.374 0.387 0.382 0.391 0.426 0.441
336 0.386 0.402 0.404 0.402 0.426 0.420 0.394 0.402 0.417 0.422 0.405 0.412 0.403 0.411 0.399 0.410 0.532 0.515 0.410 0.411 0.415 0.415 0.445 0.459
720 0.452 0.437 0.463 0.437 0.491 0.459 0.460 0.442 0.481 0.458 0.466 0.447 0.461 0.442 0.454 0.439 0.666 0.589 0.478 0.450 0.473 0.451 0.543 0.490
Avg. 0.377 0.393 0.390 0.393 0.407 0.410 0.386 0.397 0.398 0.411 0.393 0.403 0.393 0.404 0.387 0.400 0.513 0.496 0.400 0.406 0.404 0.408 0.448 0.452
ETTm2 96 0.169 0.255 0.174 0.255 0.180 0.264 0.176 0.257 0.177 0.262 0.180 0.261 0.176 0.266 0.175 0.259 0.287 0.366 0.187 0.267 0.193 0.293 0.203 0.287
192 0.235 0.299 0.243 0.302 0.250 0.309 0.243 0.303 0.247 0.307 0.246 0.306 0.240 0.307 0.241 0.302 0.414 0.492 0.249 0.309 0.284 0.361 0.269 0.328
336 0.293 0.336 0.304 0.341 0.311 0.348 0.303 0.341 0.312 0.346 0.319 0.352 0.304 0.345 0.305 0.343 0.597 0.542 0.321 0.351 0.382 0.429 0.325 0.366
720 0.390 0.393 0.399 0.397 0.412 0.407 0.400 0.398 0.414 0.403 0.405 0.401 0.406 0.400 0.402 0.400 1.730 1.042 0.408 0.403 0.558 0.525 0.421 0.415
Avg. 0.272 0.321 0.280 0.324 0.288 0.332 0.281 0.325 0.288 0.330 0.287 0.330 0.282 0.330 0.281 0.326 0.757 0.610 0.291 0.333 0.354 0.402 0.305 0.349
ETTh1 96 0.370 0.394 0.377 0.393 0.386 0.405 0.377 0.394 0.390 0.411 0.381 0.399 0.382 0.398 0.414 0.419 0.423 0.448 0.384 0.402 0.397 0.412 0.376 0.419
192 0.413 0.420 0.429 0.425 0.441 0.436 0.424 0.422 0.442 0.442 0.435 0.431 0.427 0.425 0.460 0.445 0.471 0.474 0.436 0.429 0.446 0.441 0.420 0.448
336 0.450 0.440 0.471 0.446 0.487 0.458 0.459 0.442 0.480 0.468 0.480 0.452 0.465 0.445 0.501 0.466 0.570 0.546 0.491 0.469 0.489 0.467 0.459 0.465
720 0.448 0.457 0.496 0.480 0.503 0.491 0.463 0.459 0.494 0.488 0.499 0.488 0.472 0.468 0.500 0.488 0.653 0.621 0.521 0.500 0.513 0.510 0.506 0.507
Avg. 0.420 0.428 0.443 0.436 0.454 0.447 0.431 0.429 0.452 0.452 0.449 0.442 0.437 0.434 0.469 0.454 0.529 0.522 0.458 0.450 0.461 0.457 0.440 0.460
ETTh2 96 0.283 0.337 0.296 0.345 0.297 0.349 0.292 0.343 0.328 0.371 0.297 0.347 0.309 0.359 0.302 0.348 0.745 0.584 0.340 0.374 0.340 0.394 0.358 0.397
192 0.362 0.392 0.368 0.389 0.380 0.400 0.367 0.389 0.402 0.414 0.373 0.394 0.390 0.406 0.388 0.400 0.877 0.656 0.402 0.414 0.482 0.479 0.429 0.439
336 0.404 0.424 0.411 0.422 0.428 0.432 0.412 0.424 0.435 0.443 0.410 0.426 0.426 0.444 0.426 0.433 1.043 0.731 0.452 0.452 0.591 0.541 0.496 0.487
720 0.407 0.433 0.412 0.434 0.427 0.445 0.419 0.438 0.417 0.441 0.411 0.433 0.445 0.464 0.431 0.446 1.104 0.763 0.462 0.468 0.839 0.661 0.463 0.474
Avg. 0.364 0.397 0.372 0.397 0.383 0.407 0.373 0.399 0.396 0.417 0.385 0.408 0.373 0.400 0.387 0.407 0.942 0.684 0.414 0.427 0.563 0.519 0.437 0.449
Weather 96 0.153 0.199 0.163 0.202 0.174 0.214 0.156 0.202 0.163 0.212 0.166 0.208 0.159 0.218 0.177 0.218 0.158 0.230 0.172 0.220 0.195 0.252 0.217 0.296
192 0.202 0.246 0.218 0.252 0.221 0.254 0.207 0.250 0.212 0.254 0.217 0.253 0.211 0.266 0.225 0.259 0.206 0.277 0.219 0.261 0.237 0.295 0.276 0.336
336 0.260 0.289 0.274 0.294 0.278 0.296 0.262 0.291 0.272 0.299 0.282 0.300 0.267 0.310 0.278 0.297 0.272 0.335 0.280 0.306 0.282 0.331 0.339 0.380
720 0.342 0.341 0.349 0.343 0.358 0.349 0.343 0.343 0.350 0.348 0.356 0.351 0.352 0.362 0.354 0.348 0.398 0.418 0.365 0.359 0.345 0.382 0.403 0.428
Avg. 0.239 0.269 0.251 0.273 0.258 0.279 0.242 0.272 0.249 0.278 0.255 0.278 0.247 0.289 0.259 0.281 0.259 0.315 0.259 0.287 0.265 0.315 0.309 0.360
Electricity 96 0.133 0.230 0.145 0.233 0.148 0.240 0.141 0.235 0.165 0.274 0.143 0.233 0.173 0.275 0.195 0.285 0.219 0.314 0.168 0.272 0.210 0.302 0.193 0.308
192 0.154 0.248 0.163 0.248 0.162 0.253 0.159 0.252 0.184 0.292 0.158 0.248 0.195 0.288 0.199 0.289 0.231 0.322 0.184 0.289 0.210 0.305 0.201 0.315
336 0.162 0.261 0,175 0.262 0.178 0.269 0.173 0.268 0.195 0.302 0.178 0.269 0.206 0.300 0.215 0.305 0.246 0.337 0.198 0.300 0.223 0.319 0.214 0.329
720 0.184 0.284 0.204 0.291 0.225 0.317 0.201 0.295 0.231 0.332 0.218 0.305 0.231 0.335 0.256 0.337 0.280 0.363 0.220 0.320 0.258 0.350 0.246 0.355
Avg. 0.158 0.256 0.172 0.258 0.178 0.270 0.169 0.263 0.194 0.300 0.174 0.264 0.201 0.300 0.216 0.304 0.244 0.334 0.193 0.295 0.225 0.319 0.214 0.327
Traffic 96 0.375 0.251 0.407 0.252 0.395 0.268 0.426 0.276 0.605 0.344 0.376 0.251 0.570 0.310 0.544 0.359 0.522 0.290 0.593 0.321 0.650 0.396 0.587 0.366
192 0.395 0.262 0.431 0.262 0.417 0.276 0.458 0.289 0.613 0.359 0.398 0.261 0.577 0.321 0.540 0.354 0.530 0.293 0.617 0.336 0.598 0.370 0.604 0.373
336 0.414 0.271 0.456 0.269 0.433 0.283 0.486 0.297 0.642 0.376 0.415 0.269 0.588 0.324 0.551 0.358 0.558 0.305 0.629 0.336 0.605 0.373 0.621 0.383
720 0.445 0.289 0.509 0.292 0.467 0.302 0.498 0.313 0.702 0.401 0.447 0.287 0.597 0.337 0.586 0.375 0.589 0.328 0.640 0.350 0.645 0.394 0.626 0.382
Avg. 0.407 0.268 0.451 0.269 0.428 0.282 0.467 0.294 0.641 0.370 0.409 0.267 0.583 0.323 0.555 0.362 0.550 0.304 0.620 0.336 0.625 0.383 0.610 0.376
Solar-Energy 96 0.193 0.223 0.200 0.207 0.203 0.237 0.197 0.241 0.208 0.243 0.200 0.230 0.283 0.353 0.234 0.286 0.310 0.331 0.250 0.292 0.290 0.378 0.242 0.342
192 0.226 0.249 0.228 0.233 0.233 0.261 0.231 0.263 0.258 0.281 0.229 0.253 0.316 0.374 0.267 0.310 0.734 0.725 0.296 0.318 0.320 0.398 0.285 0.380
336 0.235 0.261 0.262 0.244 0.248 0.273 0.241 0.268 0.293 0.311 0.243 0.269 0.347 0.393 0.290 0.315 0.750 0.735 0.319 0.330 0.353 0.415 0.282 0.376
720 0.239 0.268 0.258 0.249 0.249 0.275 0.250 0.281 0.290 0.315 0.245 0.272 0.348 0.389 0.289 0.317 0.769 0.765 0.338 0.337 0.357 0.413 0.357 0.427
Avg. 0.223 0.250 0.237 0.233 0.233 0.262 0.230 0.264 0.262 0.288 0.229 0.256 0.324 0.377 0.270 0.307 0.641 0.639 0.301 0.319 0.330 0.401 0.292 0.381
@@ B.3 Short-term Forecasting
Table 10 present the full results for short-term forecasting. The look-back horizon $L$ is set to $96$ and the forecasting horizon $T\in\{12,24,48\}$. Across all four PEMS datasets, TimeFilter consistently delivers the highest performance, particularly with a significant improvement on PEMS08, showcasing its effectiveness and reliability.
Models TimeFilter DUET iTransformer Leddam SOFTS PatchTST Crossformer TimesNet TiDE DLinear SCINet FEDformer
@@
MSE
MAE
PEMS03 12 0.063 0.165 0.064 0.166 0.071 0.174 0.068 0.174 0.064 0.165 0.099 0.216 0.090 0.203 0.085 0.192 0.178 0.305 0.122 0.243 0.066 0.172 0.126 0.251
24 0.079 0.185 0.081 0.186 0.093 0.201 0.094 0.202 0.083 0.188 0.142 0.259 0.121 0.240 0.118 0.223 0.257 0.371 0.201 0.317 0.085 0.198 0.149 0.275
48 0.110 0.222 0.114 0.222 0.125 0.236 0.140 0.254 0.114 0.223 0.211 0.319 0.202 0.317 0.155 0.260 0.379 0.463 0.333 0.425 0.127 0.238 0.227 0.348
Avg. 0.084 0.191 0.086 0.192 0.096 0.204 0.101 0.210 0.087 0.192 0.151 0.265 0.138 0.253 0.119 0.271 0.271 0.380 0.219 0.295 0.093 0.203 0.167 0.291
PEMS04 12 0.068 0.167 0.079 0.181 0.078 0.183 0.076 0.182 0.074 0.176 0.105 0.224 0.098 0.218 0.087 0.195 0.219 0.340 0.148 0.272 0.073 0.177 0.138 0.262
24 0.080 0.183 0.096 0.203 0.095 0.205 0.097 0.209 0.088 0.194 0.153 0.257 0.131 0.256 0.103 0.215 0.292 0.398 0.224 0.340 0.084 0.193 0.177 0.293
48 0.101 0.209 0.114 0.226 0.120 0.233 0.132 0.249 0.110 0.219 0.229 0.339 0.205 0.326 0.136 0.250 0.409 0.478 0.335 0.437 0.099 0.211 0.270 0.368
Avg. 0.083 0.186 0.096 0.203 0.098 0.207 0.102 0.213 0.091 0.196 0.162 0.273 0.145 0.267 0.109 0.220 0.307 0.405 0.236 0.350 0.085 0.194 0.195 0.308
PEMS07 12 0.055 0.150 0.060 0.156 0.067 0.165 0.066 0.164 0.057 0.152 0.095 0.207 0.094 0.200 0.082 0.181 0.173 0.304 0.115 0.242 0.068 0.171 0.109 0.225
24 0.068 0.166 0.073 0.172 0.088 0.190 0.079 0.185 0.073 0.173 0.150 0.262 0.139 0.247 0.101 0.204 0.271 0.383 0.210 0.329 0.119 0.225 0.125 0.244
48 0.089 0.193 0.096 0.201 0.110 0.215 0.115 0.228 0.096 0.195 0.253 0.340 0.311 0.369 0.134 0.238 0.446 0.495 0.398 0.458 0.149 0.237 0.165 0.288
Avg. 0.071 0.170 0.076 0.176 0.088 0.190 0.087 0.192 0.075 0.173 0.166 0.270 0.181 0.272 0.106 0.208 0.297 0.394 0.241 0.343 0.112 0.211 0.133 0.282
PEMS08 12 0.064 0.162 0.072 0.168 0.079 0.182 0.070 0.173 0.074 0.171 0.168 0.232 0.165 0.214 0.112 0.212 0.227 0.343 0.154 0.276 0.087 0.184 0.173 0.273
24 0.079 0.182 0.093 0.191 0.115 0.219 0.091 0.200 0.104 0.201 0.224 0.281 0.215 0.260 0.141 0.238 0.318 0.409 0.248 0.353 0.122 0.221 0.210 0.310
48 0.105 0.214 0.123 0.217 0.186 0.235 0.145 0.261 0.164 0.253 0.321 0.354 0.315 0.335 0.198 0.283 0.497 0.510 0.440 0.470 0.189 0.270 0.320 0.394
Avg. 0.083 0.186 0.096 0.192 0.127 0.212 0.102 0.211 0.114 0.208 0.238 0.289 0.232 0.270 0.150 0.244 0.347 0.421 0.281 0.366 0.133 0.225 0.234 0.326